### v1

# AV-Master: Dual-Path Comprehensive Perception Makes Better Audio-Visual Question Answering

Abstract
Audio-Visual Question Answering (AVQA) requires models to effectively utilize both visual and auditory modalities to answer complex and diverse questions about audio-visual scenes. However, existing methods lack sufficient flexibility and dynamic adaptability in temporal sampling and modality preference awareness, making it difficult to focus on key information based on the question. This limits their reasoning capability in complex scenarios. To address these challenges, we propose a novel framework named AV-Master. It enhances the model’s ability to extract key information from complex audio-visual scenes with substantial redundant content by dynamically modeling both temporal and modality dimensions. In the temporal dimension, we introduce a dynamic adaptive focus sampling mechanism that progressively focuses on audio-visual segments most relevant to the question, effectively mitigating redundancy and segment fragmentation in traditional sampling methods. In the modality dimension, we propose a preference-aware strategy that models each modality’s contribution independently, enabling selective activation of critical features. Furthermore, we introduce a dual-path contrastive loss to reinforce consistency and complementarity across temporal and modality dimensions, guiding the model to learn question-specific cross-modal collaborative representations. Experiments on four large-scale benchmarks show that AV-Master significantly outperforms existing methods, especially in complex reasoning tasks.
I Introduction
Humans perceive the world through various modalities, such as vision, hearing, and touch. Inspired by such multisensory experiences, researchers have increasingly focused on a range of multimodal understanding tasks [1, 2, 3]. Among these tasks, Audio-Visual Question Answering (AVQA) [4, 5, 6] is a practical task with broad application prospects. AVQA utilizes both visual and auditory modalities and requires the model to discover the associations between them to answer various questions. This process involves dynamically understanding audio-visual segments and addressing question-specific modality preferences, significantly increasing the complexity of the task.
For AVQA, we argue it is essential to focus on the following: (i) How can the model identify the most relevant visual and audio segments of the video related to the given question? (ii) How can the model extract cues from the most relevant modality information related to the given question? For (i), most current models employ discrete [7, 5] or continuous [6] sampling methods. Although these methods have achieved promising results, they still have notable limitations. The former compromises the inherent temporal cues in audio-visual segments, resulting in hallucinations during fine-grained understanding. The latter partially alleviates this issue but introduces significant redundancy, which hinders further enhancement of model performance. For (ii), most mainstream methods are not specifically designed to differentiate between various input modalities but instead treat them merely as supplementary information. A small number of researchers [8] have recognized these challenges and opted to design fusion weights for features from different modalities during the multimodal interaction process, thereby biasing the model towards a particular modality. However, this method overlooks the challenges in learning significant differences among fine-grained features and requires the model to reinterpret modality preferences in complex multimodal semantics, without explicitly enhancing the learning process, thus providing limited guidance for improving the model’s performance.
To solve these issues, we propose AV-Master, which can find the important fine-grained areas related to the current question within complex dynamic audio-visual scenes, and generate the optimal answer by integrating global preference information. As shown in Fig. 1, unlike previous work [5, 6], AV-Master employs dynamic adaptive focus sampling to extract the fine-grained focus features from the audio-visual segments relevant to the current questioning scenario. The focus sampling combines the advantages of discrete (, uniform [7], top-k [5]) and continuous (, gaussian, gaussian-experts [6]) sampling methods, allowing the model to capture all continuous time steps while significantly reducing redundant information obtained from sampling. Furthermore, we propose a global preference activation strategy focused on model modality preferences. This strategy determines the preference distribution of the model in different modalities through independent perception and improves the decision-making capabilities of the model. Compared to using dynamic weights at the multimodal fusion stage, our enhancement strategy focuses on the global information in the initial features. This avoids biases caused by decoding from complex multimodal features, while also complementing the fine-grained information obtained through temporal dynamic perception, achieving comprehensive audio-visual understanding. Our contributions are summarized as follows:
We propose a dual-path audio-visual learning model named AV-Master that enhances the understanding of audio-visual scenes by perceiving fine-grained details and global modality preferences related to questions, achieving cross-modal mapping from audio-visual signals to textual answers.
We introduce a dynamic adaptive focus sampling method that progressively performs adaptive learning from the input audio-visual segments during the encoding, enabling precise capture of the focal areas within a large number of redundant segments.
Extensive evaluation on four benchmark AVQA datasets demonstrates that our proposed AV-Master is superior and achieves new state-of-the-art performance compared to existing AVQA methods.
The remainder of this paper is organized as follows. In Section II, we review the related work relevant to the research direction of this study. Section III provides a detailed description of the architecture of the proposed model. Section IV presents experimental results that validate the effectiveness of our proposed methods. Finally, in Section V, we conclude by summarizing the main contributions of this paper.
II Related Work
II-A Audio-Visual Scene Understanding
In recent years, audio-visual scene understanding has emerged as a significant research area, garnering ever-increasing attention from the academic community. The core idea of this field originates from the human instinct to perceive the world through the synergy of multiple senses. Specifically, visual and auditory information are not only complementary but also often inseparable in understanding complex, dynamic environments. These two modalities are tightly linked through semantic consistency (e.g., the image of a dog matching the sound of its bark) and spatio-temporal correlation (a sound synchronizing with and emanating from its visual source). This interplay provides the key elements for achieving a sophisticated understanding of scenes that surpasses the capabilities of single-modality perception. Building on this foundation, researchers have explored numerous specific sub-tasks such as sound source localization [9, 10, 11], action recognition [12, 13], event detection [14, 15, 16], video parsing [17, 18, 19], audio-visual source separation [20, 21], etc. These tasks aim to fully leverage the interaction and fusion of audio and visual information to overcome the limitations of unimodal perception, thereby enhancing the fine-grained understanding of dynamic audio-visual scenes.
Within this broad research landscape, our work focuses on the more advanced cognitive task of audio-visual question answering. We achieve precise scene understanding and reasoning based on multimodal information by specifically emphasizing the temporal dynamic perception of audio-visual segments and employing a global preference activation strategy to dynamically capture the dependency of different questions on specific modalities.
II-B Audio-Visual Question Answering
The audio-visual question answering (AVQA) aims at achieving fine-grained comprehension and reasoning of complex audio-visual scenes. This task involves a comprehensive analysis of audio, visual, and their fused information, allowing the model to provide accurate answers to different questions. Existing research focuses mainly on unimodal audio question answering (AQA) and visual question answering (VQA), but a single modality is insufficient to fully capture the rich semantic information contained in natural videos. To this end, recent works [5, 8, 6] have begun exploring the field of AVQA. Among these works, Li et al. constructed declarative sentence prompts based on question templates to help the temporal awareness module better identify key segments relevant to the question, and designed a novel spatial awareness module to facilitate efficient fusion of visual tokens. Zhao et al. aligned audio-visual cues across spatial and temporal dimensions through contrastive learning, and adaptively assigned fusion weights to the visual and audio modalities according to the question. Kim et al. proposed the QA-TIGER framework, which adaptively focuses on both continuous and discontinuous frames according to the question, explicitly injects question information, and applies progressive optimization.
However, existing works only adopt simple audio-visual segment selection methods, which introduce a large amount of redundant information, making it difficult for the subsequent decoder to extract key clues from the coarse-grained multimodal features. Moreover, they overlook the fact that different questions may have varying demands on visual and audio modalities. As a result, the models often fail to dynamically leverage the relevant dominant modality for different questions and instead perform a coarse multimodal fusion, which, to some extent, undermines the overall performance of the model. In contrast, our work enables fine-grained key audio-visual learning through dynamic adaptive sampling based on the question, and enhances the model from a global perspective by leveraging modality preference, ultimately leading to more accurate answers.
III Methodology
In this section, we will introduce our proposed AV-Master in detail. Specifically, we first introduce the feature information used and the feature extraction settings in Section III-A. Subsequently, in Section III-B and Section III-C, we respectively explain the proposed temporal dynamic perception path and global preference activation path. Finally, in Section III-D, we describe in detail the learning objectives involved in the training phase, which include the dual-path prediction loss and the dual-path contrastive loss. The overall architecture is shown in Fig. 2, and the specific implementation flow of the temporal dynamic perception path proposed in Fig. 2 is illustrated in Fig. 3.
III-A Input Representation
(a) Visual representation: For a given video, we split it into non-overlapping 1 segments, each with paired audio and visual elements. Each visual segment is processed by a pre-trained vision-language model CLIP [22]. In this process, a special token is added at the beginning of each segment and is used as the visual feature. The visual features can be represented as , where denotes the feature dimension.
(b) Audio representation: For each audio segment, we follow previous works [5, 4, 6] that use the pre-trained VGGish model [23] to extract audio features. The audio features can be represented as . The parameters of the CLIP and VGGish models are frozen during training.
(c) Question representation: For the input question, we use the CLIP text encoder to obtain word-level features , and extract the sentence-level feature by taking the embedding of the first token. denotes the number of question tokens.
represent the input predefined CLS tokens (serve as audio-visual templates),
represent the audio-visual features at a certain time step, and
is a learned bias.III-B Temporal Dynamic Perception Path
To extract key information from complex initial audio-visual features, we propose a temporal dynamic perception path, which consists of two modules: an audio-visual focus capture module and an audio-visual key fusion module. The former focuses more on the relationships within each modality, aiming to discover key hidden clues in the current modality, while the latter concentrates on establishing connections between different modalities. Additionally, the inputs to the perception path are the outputs from the previous feature encoding stage, including visual features , audio features , word-level features , and sentence-level features .
(a) Audio-visual focus capture: As shown in Fig. 3 (a), to achieve fine-grained perception, the visual features and the audio features are fed into the audio-visual focus capture module for dynamic adaptive focus sampling at each time step from time to time , where . During the perception process, we utilize predefined templates ( and ) to focus on the audio-visual features at each moment, thereby extracting global effective information from the complex initial features. Moreover, we introduce learnable biases to highlight key regions with significant variations between different moments for improving sampling accuracy. Over time, the predefined templates progressively extract the critical information hidden at different moments throughout the entire audio-visual segment. Finally, we take the templates from the -th time step as the output and obtain the fine-grained focus features and .
(b) Dynamic adaptive focus sampling: For a specific time step , the detailed focus sampling process is shown in Fig. 3 (b). First, the -th time step visual feature and audio feature together with the -th time step template and serve as inputs for the entire sampling process. Features and are repeated to match the length of their corresponding templates and then added to the templates. The resulting sums are fed into a self-attention block (SAB) for attention enhancement. These enhanced features are then added to a learnable bias and go through the above operation once more. Afterward, these features, along with the repeated feature, are passed into a cross-attention block (CAB) for modality interaction. Finally, we can obtain the updated templates and for the current time step. This process is formulated as follows:
where [] and represents the total number of video 1s segments. The focus sampling is formulated as:
where serves as and serves as in . For the -th time step audio feature , the same applies following the above Eqs. (1-2).
(c) Audio-visual key fusion: After obtaining the fine-grained audio-visual focal features and , we feed them into the audio-visual key fusion module, where multimodal fusion is performed under the guidance of the word-level question feature to provide a refined semantic anchor for the subsequent decoding. The detailed audio-visual key fusion process is shown in Fig. 3 (c), and are transformed by linear layers and then added to the multimodal feature . Subsequently, each of these is fed into an SAB for enhancement, followed by a pooling operation, and then concatenated with along the feature dimension. Finally, the result is passed through a linear layer and max sampling to obtain the final fused feature . Where, is formed by concatenating the visual feature , audio feature , and question feature along the sequence length. The fusion process is formulated as follows:
where the performs summation along the sequence dimension, while the takes the maximum value along the feature dimension. The fused feature and represents the intermediate outputs.
III-C Global Preference Activation Path
Considering the potential mismatch between audio-visual segments and the possibility that the current question may be more biased toward a specific modality or scene in audio-visual question answering, we propose a global preference activation path. The preference activation path is designed to independently perceive and decouple auditory and visual inputs, activating the global contextual information within them. It serves to complement the fine-grained features from the temporal dynamic perception path, providing the model with an additional auxiliary perspective to achieve a more comprehensive understanding of the audio-visual scene.
As shown in Fig 2 (b), the visual feature and the word-level question feature are first fed into a cross-attention block. Meanwhile, the is also separately enhanced via a self-attention block. The outputs are then summed and passed through a multi-layer perceptron (MLP) to obtain the visual preference feature . Similarly, the audio feature goes through the same procedure to obtain the audio preference feature . The calculation process is as follows:
where visual feature or audio feature serves as and serves as in . The activated preference features and .
III-D Optimization and Answer Prediction
During training, the fused feature and the sentence-level question feature are jointly fed into the multimodal decoder to generate the answer prediction distribution, which is then used together with the ground-truth labels to compute the answer loss (denoted as ). In the preference activation path, features and are separately input into two independent audio/visual decoders, which also generate prediction distributions under the guidance of , resulting in two preference losses (denoted as and ). Moreover, a contrastive loss (denoted as ) is applied between the dynamic perception and preference activation paths, aiming to enhance the stability of the dual-path paradigm and improve the model’s discriminative ability by leveraging hard negative samples. During inference, we sum all the predicted distributions from the model and use the argmax function to obtain the final answer.
(a) Dual-path prediction loss: For the answer loss , we follow the standard training procedure for the AVQA. The goal of the model is to minimize the negative log-likelihood of the probabilities over multiple answer choices generated by the multimodal decoder. For the preference losses and , we also adopt a similar computation method to calculate them based on the probability distributions generated by the two audio/visual decoders. The formula for the above loss calculation can be represented as:
where is the total number of answer choices and is the set of learnable parameters of the decoders. Both the multimodal decoder and the audio/visual decoders include transformer blocks and linear layers.
Dataset # Videos # Train QA # Valid QA # Test QA AVQA 57,015 40,425 - 16,910 MUSIC-AVQA 9,288 31,904 4,568 9,129 MUSIC-AVQA-R 9,288 - - 211,572 MUSIC-AVQA-v2.0 10,492 37,408 5,346 10,819
(b) Dual-path contrastive loss: For the contrastive loss , we enhance the stability of the dual-path architecture and improve the accuracy of joint prediction by increasing the similarity between the positive sample feature from the dynamic perception path and from the preference activation path, . Meanwhile, we compare the feature from the positive sample with the feature from the negative sample and reduce their similarity to enhance the model’s ability to distinguish positive samples. The contrastive is expressed as follows:
where is the number of negative pairs and is the cosine function used to compute similarity. represents the operation of averaging the feature along the sequence length dimension, denotes the temperature. The overall loss is the weighted sum of the above losses:
where , , , and are hyperparameters used to trade-off each loss functions.
[…]
IV Experiments
IV-A Experimental Setting
(a) Dataset and evaluation metric: In this paper, we validate our proposal on four datasets: MUSIC-AVQA [7], MUSIC-AVQA-R [24], MUSIC-AVQA-v2.0 [25], and AVQA [26]. The details are summarized in Table I. Consistent with previous work, the ablation and analysis experiments are conducted on the MUSIC-AVQA dataset by default.
MUSIC-AVQA [7] is a large-scale AVQA dataset focused on multimodal understanding and reasoning in music performance scenarios. It contains 45,601 question-answer pairs distributed across 9,288 videos, with a total duration of over 150 hours. The videos are primarily collected from YouTube and cover 22 types of musical instruments (such as guitar, cello, marimba, etc.). Annotations are manually created by human annotators. The question-answer pairs are divided into three modality scenarios (audio, visual, and audio-visual), encompassing nine question types (such as existential, location, counting, etc.) and 33 question templates. The answer set includes 42 different answers. The questions require fine-grained scene understanding and spatiotemporal reasoning of both audio and visual content, for example: “Is a certain instrument present?” or “Where is the source of the sound located?”. The MUSIC-AVQA dataset emphasizes the interaction between audio and visual modalities in music performance scenes. Compared to other video question answering datasets (such as general video datasets), it places a stronger focus on audio-visual correlation.
MUSIC-AVQA-R [24] is an extended version of the MUSIC-AVQA, designed to evaluate the robustness of AVQA models and to address the issues of limited question expression variety and potential biases in the original dataset. It is constructed by rephrasing and splitting the questions in the MUSIC-AVQA test set. Specifically, researchers used AI tools to rephrase the 9,129 questions in the test set 25 times, generating semantically equivalent but more diverse expressions. Then, three annotators independently voted to retain most of the questions. As a result, the number of test questions expanded from 9,129 to 211,572, and the vocabulary size increased from 93 to 465. All questions are divided into two subsets: frequent (head) and rare (tail), classified based on the frequency of answer occurrences. Compared to MUSIC-AVQA, MUSIC-AVQA-R enhances question diversity and is better suited for testing AVQA models under conditions of diverse questions and imbalanced answer distributions.
MUSIC-AVQA-v2.0 [25] is an improved version of the MUSIC-AVQA, aimed at addressing the data bias issues in the original dataset and constructing a more balanced and challenging dataset. Researchers manually collected 1,230 new instrument performance videos (sourced from YouTube) and created 8,100 new question-answer pairs to supplement the original MUSIC-AVQA dataset. They ensured a more even answer distribution across each question category and subcategory, especially for binary questions, where the answers are nearly evenly distributed to avoid significant bias. Some of the videos were horizontally flipped to generate symmetric question-answer pairs, enhancing data diversity. MUSIC-AVQA-v2.0 resolves the answer bias problems present in the MUSIC-AVQA dataset, making it a more reliable benchmark suitable for testing model performance in unbiased AVQA tasks.
To evaluate the performance of AV-Master in scenarios beyond musical performance, we additionally introduce the AVQA dataset. AVQA [26] is a large-scale benchmark specifically designed for audio-visual question answering in real-world settings. The dataset contains 57,015 real-life videos and 57,335 question-answer pairs, with a total duration of over 158 hours. The videos are sourced from the VGG-Sound dataset [27] and cover 165 categories of daily activities and natural sounds. The questions are manually designed to ensure reliance on both modalities for reasoning, involving various types of relations such as existence, location, temporal, causal, and intentional. With its large scale, diversity, and high-quality human annotations, the dataset has become a widely used benchmark for evaluating multimodal fusion methods in complex real-world scenarios.
[…]
(b) Implementation details: For a fair comparison, we follow previous work [6] and adopt a similar setup: videos are uniformly sampled at a rate of 1 frame per second. Audio representations are extracted using the pre-trained VGGish model [23], while visual inputs and corresponding questions are encoded through the CLIP-ViT-L/14 model [22]. All extracted features are projected into a 512-dimensional space via a linear transformation. The predefined templates, and , are both composed of a set of learnable embeddings, each with a length of 8 and randomly initialized. The model is optimized using Adam [28] and starts with a learning rate of 1e-4, which is reduced by a factor of 0.1 every 8 epochs. The batch size is set to 32, and the model is trained for 30 epochs. Our proposed model is trained on NVIDIA GeForce RTX 4090 and implemented in PyTorch.
IV-B Quantitative Results
# Methods Average Accuracy (%) A-QA V-QA AV-QA Avg 1 w/o. ALL 73.37 79.23 69.82 72.94 2 w/o. AVFC 79.83 85.55 74.04 78.11 3 w/o. DPCL 78.83 85.71 73.78 77.84 4 w/o. GPAP 77.90 85.05 73.10 77.12 5 w/o. TDPP 78.77 83.73 72.94 76.83 6 AV-Master 79.95 86.58 74.22 78.51
# Average Accuracy (%) A-QA V-QA AV-QA Avg 1 ✓ 77.84 84.31 71.11 75.80 2 ✓ ✓ 78.15 84.43 70.76 75.69 3 ✓ ✓ ✓ 79.83 82.74 72.49 76.50 4 ✓ ✓ ✓ 78.83 86.33 72.63 77.36 5 ✓ ✓ ✓ 78.83 85.71 73.78 77.84 6 ✓ ✓ ✓ ✓ 79.95 86.58 74.22 78.51
(a) Compared methods: To evaluate the effectiveness and superiority of our model, we compare AV-Master with existing state-of-the-art AVQA methods across multiple datasets. These methods include: QA-TIGER [6], CoQo [29], SHMamba [30], AVAF-Net [8], PSOT [31], SaSR-Net [32], PSTP-Net [4], TSPM [5], APL [33], LAST [25], MCD [34], QAGL [35], LAVISH [36], ST-AVQA [7], Pano-AVQA [37], ACRTransformer [38], HGA [39], HCRN [40], FCNLSTM [41], LADNet [42], AVSD [43], HME [44], PSAC [45], MCAN [46] and HCAttn [47].
(b) MUSIC-AVQA: As shown in Table II, AV-Master achieves an overall accuracy of 78.51%, outperforming all existing models, including the state-of-the-art method QA-TIGER (77.62%). Notably, our model demonstrates strong performance on complex reasoning tasks such as counting, significantly outperforming the second-best method (i.e., A-Counting: 87.02% vs. 84.86%, V-Counting: 86.55% vs. 83.96%, AV-Counting: 79.13% vs. 78.58%).
(c) MUSIC-AVQA-R: As shown in Table III, compared to the MUSIC-AVQA dataset, AV-Master demonstrates more significant improvements (+3.20%) on the MUSIC-AVQA-R dataset, achieving an overall accuracy of 71.19% and setting a new state-of-the-art performance. The notable performance gain can be attributed to AV-Master’s dual-path learning paradigm, which provides the model with powerful generalization capabilities.
(d) MUSIC-AVQA-v2.0: As shown in Table IV, AV-Master outperforms existing models across all types. Notably, when trained on the biased dataset, AV-Master still achieves significant improvements over the second-best method on both the balanced and biased test sets. These results highlight the robustness and adaptability of AV-Master in handling various training environments, demonstrating its strong generalization capability even under distributional bias.
(e) AVQA: To further validate the generalization capability of our model in real-world scenarios, we conducted experiments on the AVQA dataset. As shown in Tab. V, AV-Master achieved an overall accuracy of 91.4%, surpassing all previous methods both with and without the HAVF [26] module. These results strongly demonstrate that AV-Master maintains its exceptional performance in complex, real-world audio-visual question answering tasks. It is worth noting that although the performance improvement of AV-Master on the AVQA dataset may seem limited compared to its performance on the MUSIC-AVQA series of datasets, this is primarily due to the shorter duration and simpler audio-visual content of the AVQA dataset.
(f) Comparison with pretraining-based methods: In this work, our approach focuses on designing efficient AVQA expert models that achieve competitive performance under limited training data and hardware conditions by developing powerful modules. This is also the mainstream direction in current AVQA research [6, 29, 32]. Additionally, there are some methods [48, 49, 50] that rely on large-scale pretraining, which attempt to apply large language models to audio-visual scenarios to handle downstream AVQA tasks. However, pretraining-based AVQA methods require substantial computational resources and multimodal data, and their generalization ability in specific audio-visual scenarios falls short of expectations. We present relevant experiments to compare AV-Master with pretraining-based models, including OneLLM [51], Chatbridge [52], CAT [53], CAT+ [54], VideoLLaMa [55], AVLLM [56], AVicuna [57], VideoLLaMa2 [48], CAD [58], VAST [49] and VALOR [50].
As shown in Table VI, AV-Master achieves competitive performance (78.5% accuracy) compared to various pretraining-based models, despite having significantly fewer parameters (61M) and not relying on large-scale pretraining. In contrast, methods such as CAT+ and AVicuna, which are based on large language models, achieve notably lower performance (50.1% and 49.6% accuracy) under zero-shot conditions. Although these approaches leverage abundant resources and powerful backbone models, they often exhibit weaker generalization capabilities in domain-specific audio-visual tasks. Other pretraining-based models achieve impressive performance (80.7% accuracy) after fine-tuning on the MUSIC-AVQA dataset. However, our method achieves comparable results using significantly less training data, fewer trainable parameters, and a more lightweight audio-visual feature encoder. This highlights the practicality of AV-Master and its suitability for resource-constrained audio-visual scenarios.
# Lengths Average Accuracy (%) A-QA V-QA AV-QA Avg 1 16 80.51 85.96 74.04 78.34 2 12 80.26 85.88 74.08 78.30 3 8 79.95 86.58 74.22 78.51 4 4 80.63 85.22 73.92 78.10 5 2 80.38 85.84 73.98 78.26
# A-S B-S Average Accuracy (%) A-QA V-QA AV-QA Avg 1 79.70 85.80 73.92 78.09 2 ✓ 79.95 86.58 74.22 78.51 3 ✓ 80.26 85.88 73.59 78.03 4 ✓ ✓ 80.07 86.29 73.88 78.27
Visual Encoder Audio Encoder Methods Average Accuracy (%) A-QA V-QA AV-QA Avg Resnet-18 VGGish ST-AVQA 74.06 74.00 69.54 71.52 AV-Master 78.21 79.23 70.25 74.04 CLIP VGGish PSTP-Net 70.91 77.26 72.57 73.52 AV-Master 78.40 82.25 72.82 76.31 CLIP VGGish TSPM 76.91 83.61 73.51 76.79 QA-TIGER 78.58 85.14 73.74 77.62 AV-Master 79.95 86.58 74.22 78.51 CLAP PSOT 79.08 87.12 74.07 78.42 AV-Master 80.63 87.78 74.92 79.34 Internvideo2 Internvideo2 CoQo 79.27 87.90 75.80 79.60 AV-Master 81.50 88.32 75.86 80.15
IV-C Ablation Studies
(a) Ablation study on main components: To explore the effectiveness of each component, we removed them individually and re-evaluated performance. As shown in Tab. VII, removing different components leads to varying degrees of performance degradation for AV-Master. Specifically, when the TDPP (Temporal Dynamic Perception Path) is removed, the performance drops to 76.83%; removing the GPAP (Global Preference Activation Path) results in a drop to 77.12%; removing the DPCL (Dual-path Contrastive Loss) lowers performance to 77.84%, and removing the AVFC (Audio-Visual Focus Capture) decreases it to 78.11%. When all components are removed simultaneously, the performance drops significantly from 78.51% to 72.94%. These results show that each component in AV-Master contributes to performance improvement, and the best results are achieved only when all components are present.
(b) Impact of template lengths: We explored the impact of the lengths of predefined templates and on model performance. As shown in Tab. IX, the average accuracy of the model reaches its highest when the template length is set to 8. If the template length is too short, important information may be missed during sampling, leading to reduced performance. Conversely, if the template length is too long, redundant information may be introduced, which can impair the model’s performance.
(c) Impact of weight sharing strategies: We also explored the impact of different weight-sharing strategies in focus sampling. As shown in Tab. X, the model achieves the best overall performance when attention blocks share weights while biases remain unshared — this configuration is also adopted as our default setting. Interestingly, when only biases are shared, the model attains the lowest overall accuracy but achieves the highest accuracy on the A-QA task. This suggests that bias sharing may benefit certain subtasks while potentially hindering overall model performance.
(d) Ablation study on loss functions: To analyze the contribution of each loss function, we conducted an ablation study with the results presented in the Tab. VIII. This study systematically evaluates the model’s performance by incrementally adding different loss components. The baseline model, trained only with the answer loss , achieves an average accuracy of 75.80%. Consecutively adding the contrastive loss and the preference losses and provides progressive gains. The final model, which integrates all four loss functions, achieves the highest average accuracy of 78.51%. This demonstrates that each loss component plays a vital role, and their combined effect is crucial for optimizing the model’s overall performance. It is worth noting that, as can be seen from the results in the third, fourth, and fifth rows of the table, visual preference training provides the most significant overall improvement to the model compared to other training objectives (aside from basic ). This also indicates that in the vast majority of scenarios, the model relies more heavily on visual information for question answering.
(e) Impact of different feature extractors: The results in Table XI demonstrate that different visual encoders have a significant impact on the performance of AV-Master. When using ResNet-18 [59] as the visual encoder, AV-Master achieves an average accuracy of 74.04%, showing a considerable improvement compared to ST-AVQA (71.52%) with the same visual encoder. When a more powerful visual encoder, CLIP [22], is adopted, the average accuracy of AV-Master further increases to 76.31%, surpassing PSTP-Net (73.52%). Moreover, when using CLIP [22] and Internvideo2 [60] as visual encoders, AV-Master achieved the best results across all subtasks (A-QA, V-QA, AV-QA), significantly outperforming comparable methods. In summary, as the visual encoder was progressively upgraded from ResNet-18 to Internvideo2, the performance of AV-Master on all tasks steadily improved.
Furthermore, the choice of audio encoders also plays a crucial role. When fixing the visual encoder to CLIP, switching from VGGish [23] to a more advanced audio encoder like CLAP [61] boosts the average accuracy from 78.51% to 79.34%, outperforming its competitor PSOT (78.42%). This trend continues with the most powerful backbones; using Internvideo2 for both visual and audio encoding, AV-Master achieves the highest overall average accuracy of 80.15%, again surpassing the competing method CoQo (79.60%). These results collectively demonstrate that the performance of AV-Master is consistently enhanced by leveraging stronger feature extractors for both the visual and audio modalities, underscoring the importance of high-quality unimodal representations for complex audio-visual reasoning tasks.
(f) Impact of different audio-visual segment lengths: To investigate the effect of audio-visual segment lengths on model performance and the overall robustness of AV-Master in scenarios with limited audio-visual input, we conducted an ablation study on the segment length. As shown in Fig. 4, when the amount of audio–visual input decreases, the overall performance of all models shows a downward trend, indicating that more complete audio-visual information can provide richer cues and is beneficial for improving model performance. Meanwhile, compared with QA-TIGER, AV-Master maintains a relatively stable performance decline when facing reduced audio-visual content, suggesting that it is capable of extracting and integrating key features from limited audio-visual information. This can be attributed to the proposed audio–visual focus capture module, which progressively refines coarse-grained audio–visual features into fine-grained cues. In addition, the involvement of the preference activation strategy and the dual-path model architecture further strengthen the model’s robustness. Note that when the segment length is 48, AV-Master’s performance shows a slight drop. This may reflect the model reaching near saturation during the middle-to-late period or being less sensitive to redundant tail segments.
# Methods Average Accuracy (%) A-QA V-QA AV-QA Avg 1 w/o. APE 79.21 86.50 73.31 77.85 2 w/o. VPE 79.83 84.02 73.51 77.41 3 w/o. AVPE 77.65 85.59 72.98 77.15 4 AV-Master 79.95 86.58 74.22 78.51
(g) Ablation study on input modalities: To investigate the contribution of different input modalities and to verify the stability of the AVQA model, we conduct an ablation study using two input settings: without visual modal (A+Q) and without audio modal (V+Q). As shown in Fig. 5, all models demonstrate improved performance when the visual modality is present. Specifically, AV-Master achieves the highest average accuracy across both input settings, reaching 69.8% with A+Q and 76.9% with V+Q. Compared to the baseline model (QA-TIGER), AV-Master shows consistent improvements across all sub-tasks (A-QA, V-QA, AV-QA), suggesting its superior ability to extract and fuse relevant features from the input. Notably, the performance gap between A+Q and V+Q settings is larger for QA-TIGER and AV-Master than for AVST, indicating that advanced models are more effective at leveraging visual cues. These results highlight the dominant role of the visual modality in multimodal question answering, while also confirming the robustness of AV-Master under varying modality conditions. The results for QA-TIGER in (f) and (g) were reproduced using the official code.
# Methods Average Accuracy (%) A-QA V-QA AV-QA Avg 1 W-ADD 78.71 81.79 71.86 75.70 2 MUL 79.70 85.30 73.70 77.84 3 ADD 79.95 86.58 74.22 78.51
(h) Ablation study on modality preference enhancement: To further investigate the effectiveness of our proposed modality preference enhancement strategy and to identify which modality has the greatest impact on overall model performance, we conduct an ablation study by selectively disabling different components during the inference stage. As shown in Table XII, removing the audio-visual preference enhancement (AVPE) leads to the most significant performance drop, reducing the average accuracy to 77.15%, which highlights the importance of jointly modeling audio-visual scenarios. Additionally, disabling visual preference enhancement (VPE) results in lower V-QA accuracy (84.02% vs. 86.58%) and overall average accuracy (77.41%), indicating that visual preference plays a critical role in AVQA. Similarly, removing the audio preference enhancement (APE) leads to degraded performance on A-QA and AV-QA, showing the necessity of modeling audio preference. Overall, the full model achieves the best performance across all metrics, confirming the complementary benefits of modality-specific preference enhancement and their joint contribution to robust AVQA performance.
(i) Impact of different enhancement methods: To explore the impact of different preference distribution fusion methods on model performance, we conduct a comparative study using three representative methods: (1) W-ADD, which first computes a weighted ratio based on the initial overall scores of the visual and auditory preference distributions before performing a weighted summation; (2) MUL, which applies element-wise multiplication between modality preferences; and (3) ADD, a simple summation of the two distributions. As shown in Table XIII, the ADD strategy achieves the best overall performance, with an average accuracy of 78.51%, outperforming both MUL (77.84%) and W-ADD (75.70%). This result suggests that a direct and unweighted summation of visual and auditory preference distributions can better preserve cross-modal complementary information and prevent overfitting to any single modality. Interestingly, although the W-ADD method introduces an adaptive weighting mechanism, its performance lags behind due to potential imbalances introduced by dynamic scaling. When viewed together with the previous ablation results in Table XII, it is evident that both the design of modality-specific preference enhancements and the fusion mechanism of these preferences play a crucial role in optimizing overall AVQA performance.
(j) Impact of different trade-off hyperparameters: As shown in Fig. 7, we investigate the impact of four trade-off hyperparameters , , and , which are used to balance different loss functions. To assess the importance of each loss function in AV-Master, we vary one hyperparameter at a time while keeping the others fixed at a default value of 1.0. The experimental results show that increasing from 0.4 to 1.0 significantly improves accuracy. However, when the value increases beyond a certain threshold (e.g., 1.6), accuracy declines slightly, suggesting that overemphasizing a single objective may diminish the contributions of others. Similar patterns are observed for , and , underscoring the independent yet equally important roles of each objective in the AV-Master. Based on these findings, we set all trade-off parameters to a default value of 1.0 to ensure a balanced contribution from each loss component. This configuration achieves optimal overall performance.
IV-D Qualitative Analysis
(a) Focus effects at different time steps: As shown in Fig. 6, we present the variation in accuracy over time during the audio-visual focus capture process. The blue dots in the figure represent the original accuracy data at different moments, while the red curve indicates the overall trend. Although the original data exhibits some fluctuations, the trend line shows that the accuracy remains relatively stable during the first 40 steps. However, starting at step 45, the accuracy begins to rise significantly, reaching its peak at step 60. This suggests that the audio-visual focusing mechanism becomes more effective in the later stages, and the model progressively improves its ability to extract and fuse relevant information during the focusing process.
(b) Attention visualization of different modalities: As shown in Fig. 8, we illustrate which parts of the text the visual and audio modalities attend to within the global preference activation path. For the input question “Where is the loudest instrument?”, the visual modality primarily attends to the word “instrument” while paying minimal attention to the audio-related word “loudest.” In contrast, the audio modality places significant focus on “loudest.” These results indicate that AV-Master can effectively distinguish between visual and auditory cues and accurately align them with the corresponding textual elements, which further demonstrates AV-Master’s capability in cross-modal understanding.
(c) Qualitative results comparison of different models: Fig. 9 presents a qualitative comparison among our proposed AV-Master, MLLM (VideoLLaMA2 [48] and Qwen3-Max [62]), and another AVQA expert model (QA-TIGER [6]) across four distinct audio-visual scenarios. In each case, AV-Master exhibits a superior ability to comprehend audio-visual content and provides accurate answers that align closely with both visual and auditory cues. In contrast, the other models fail to achieve such consistency. For example, in the first scenario (a), both VideoLLaMA2 and QA-TIGER incorrectly identify the instrument as an “accordion”, whereas AV-Master correctly identifies it as a “drum”, owing to its accurate interpretation of the audio signal. Similarly, in the other scenarios (c-d), AV-Master accurately infers the spatial and temporal characteristics of sound events, while the competing models fail to effectively integrate these multimodal cues. These qualitative results reinforce our quantitative findings, showing that AV-Master is more robust in fine-grained audio-visual reasoning, particularly in scenarios that require synchronized and cross-modal understanding. This further validates the effectiveness of our carefully designed modules in fully leveraging both audio and visual modality information. The inference results for VideoLLaMA2 and Qwen3-Max were obtained from its official demo, simulating application scenarios in low-resource environments.
V Conclusion
In this paper, we propose AV-Master, a novel dual-path audio-visual question answering expert model designed to address the challenges faced by existing models in processing complex audio-visual scenes. Current methods often struggle to flexibly focus on the most question-relevant spatiotemporal segments when confronted with a large amount of redundant information, and they lack the ability to dynamically perceive the importance of different modalities for different questions. This limits their comprehensive understanding of audio-visual scenes. To tackle these difficulties, AV-Master introduces a dynamic adaptive focus sampling mechanism and a global modality preference activation strategy. These enable it to effectively capture question-relevant audio-visual segments and modality preferences, thereby enhancing its decision-making capabilities. Extensive experiments on multiple large-scale AVQA benchmark demonstrate that by meticulously capturing key audio-visual details and integrating a global understanding of modality preferences, AV-Master provides an efficient and powerful solution for AVQA, particularly excelling in complex reasoning tasks and under challenging data distributions.